If possible answer all. Most imporant are 2 and 3

1. A company is projected to have a free cash flow of $386 million next year, growing at a 5.1% rate until the end of year 3. After that, cash flows are expected to grow at a stable rate of 2.4%. The company's cost of capital is 11.9%. The company owes $93 million to lenders and has $14 million in cash. If it has 243 million shares outstanding, what is your estimate for its stock price? Round to one decimal place.

2. A company is projected to generate free cash flows of $49 million per year for the next two years, after which it is projected grow at a steady rate in perpetuity. The company's cost of capital is 10.9%. It has $24 million worth of debt and $6 million of cash. There are 14 million shares outstanding. If the appropriate terminal exit value for this company is 14, what's your estimate of the company's stock price? Round to one decimal place.

3. You are valuing Soda City Inc. It has $121 million of debt, $81 million of cash, and 171 million shares outstanding. You estimate its cost of capital is 10.9%. You forecast that it will generate revenues of $717 million and $783 million over the next two years. Projected operating profit margin is 29%, tax rate is 26%, reinvestment rate is 37%, and terminal exit value multiple at the end of year 2 is 12. What is your estimate of its share price? Round to one decimal place. ​[Hint: Compute projected FCFF for years 1 and 2 based on info provided, compute terminal value using the exit multiple method, discount it all to find EV, walk the bridge to Equity, divide by number of shares outstanding.]

4. You are asked to value Gamecocks Inc. using the relative valuation method. Gamecocks Inc.'s earnings forecast for next year (EPS (next year)) is $2.44. The valuation and earnings of comparable companies are provided below. What is your estimate for the company's stock price? Round to one decimal place. ​[Hint: Compute the corresponding valuation multiple of the comparables and take simple average across the three comparable companies. Then apply this average multiple to Gamecocks Inc.]

stock price EPS (TTM) EPS (next year) EPS growth 5Y
A $18.3 $1.04 $1.22 5.1%
B $26.3 $1.12 $1.42 6.9%
C $55.9 $2.11 $2.80 10.2%
Answer #1

1]

Stock price = equity value / shares outstanding.

Equity value = firm value - debt + cash.

Firm value = present value of cash flows of next 3 years + present value of terminal value after 3 years.

Present value = future value / (1 + cost of capital)number of years

Terminal value after 3 years = Year 4 cash flow / (cost of capital - constant growth rate).

Year 4 cash flow = $426,375,986 * (1 + 2.4%)  

Terminal value after 3 years = $426,375,986 * (1 + 2.4%) / (11.9% - 2.4%) = $4,595,884,312.

Present value of terminal value = $4,595,884,312 / (1 + 11.9%)3 = $3,280,037,648.

Present value of next 3 years cash flows = $973,239,905.

fx =SUM(C2:04) C5 А в 1 Year Cash Flow 1 $386,000,000 2 $405,686,000 3 $426,375,986 Total Present Value $344,950,849 $323,988

А 1 Year 2 1 3 2 43 5 Cash Flow 386000000 =B2*(1+5.1%) =B3*(1+5.1%) Total C Present Value =B2/(1+11.9%) A2 =B3/(1+11.9%)^A3 l

Firm value =  $3,280,037,648 + $973,239,905 = $4,253,277,553.

Equity value = firm value - debt + cash.

Equity value =  $4,253,277,553 - $93,000,000 + $14,000,000

Equity value = $4,174,277,553.

Stock price = equity value / shares outstanding

Stock price = $4,174,277,553 / 243,000,000.

Stock price = $17.2.